OpenStudy (anonymous):

Which of the following statements best describes the graph of 4x – y = 1? It is a straight line joining the points (0, –1), (1, 3), and (–1, –5). It is a straight line joining the points (4, –3), (–1, 2), and (–4, 5). It is a curve joining the points (4, –1), (2, 3), and (4, 1). It is a curve joining the points (0, –1), (-1, –3), and (1, 5). @chosenmatt

OpenStudy (anonymous):

4x-y=1 @chosenmatt i think its a

OpenStudy (anonymous):

@itsbribro

OpenStudy (itsbribro):

yup its a

OpenStudy (anonymous):

i got one more question on my 45 question test can you help!!

OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8 -2y + 6 2y - 6 2y + 6 -2y - 6